# Lensless fiber endomicroscopic phase imaging with speckle-conditioned diffusion model

### v1

@@
Lensless fiber endomicroscopic phase imaging with speckle-conditioned diffusion model
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Lensless fiber endomicroscope is an emerging tool for in-vivo microscopic imaging, where quantitative phase imaging (QPI) can be utilized as a label-free method to enhance image contrast. However, existing single-shot phase reconstruction methods through lensless fiber endomicroscope typically perform well on simple images but struggle with complex microscopic structures. Here, we propose a speckle-conditioned diffusion model (SpecDiffusion), which reconstructs phase images directly from speckles captured at the detection side of a multi-core fiber (MCF). Unlike conventional neural networks, SpecDiffusion employs iterative phase denoising steps for speckle-driven phase reconstruction. The iteration scheme allows SpecDiffusion to break down the phase reconstruction process into multiple steps, gradually building up to the final phase image. This attribute alleviates the computation challenge at each step and enables the reconstruction of rich details in complex microscopic images. To validate its efficacy, we build an optical system to capture speckles from MCF and construct a dataset consisting of $100,000$ paired images. SpecDiffusion provides high-fidelity phase reconstruction results and shows powerful generalization capacity for unseen objects, such as test charts and biological tissues, reducing the average mean absolute error of the reconstructed tissue images by $7$ times. Furthermore, the reconstructed tissue images using SpecDiffusion shows higher accuracy in zero-shot cell segmentation tasks compared to the conventional method, demonstrating the potential for further cell morphology analysis through the learning-based lensless fiber endomicroscope. SpecDiffusion offers a precise and generalized method to phase reconstruction through scattering media, including MCFs, opening new perspective in lensless fiber endomicroscopic imaging.
1 Introduction
@@
Lensless fiber endomicroscope has been an emerging tool in medical imaging and diagnostics, noted for its high imaging resolution and minimum invasiveness [1, 2, 3, 4, 5, 6, 7, 8, 9]. Unlike traditional endoscopes that rely on lenses on the distal fiber tip to capture and transmit images, lensless fiber endomicroscope employs computational imaging techniques to visualize microscopic structures with an ultra-thin probe. Therefore, it enables microscopic exploration of areas previously inaccessible, minimizes patient discomfort and contributes to faster post-operation recovery [10]. For typical lensless fiber endomicroscopes employing wavefront shaping technologies, in-vivo confocal fluorescent imaging is usually utilized as auxiliary [11, 12, 13, 14]. However, the potential toxin of fluorescent dyes has prompted the exploration of safer alternatives [13]. Quantitative phase imaging (QPI) serves as a promising label-free technique, which utilizes the phase information to enhance endomicroscopic imaging capabilities [5]. Moreover, it enables the extraction of critical biophysical properties of biological samples, such as refractive index [15, 16], cell volume [17] and dry mass [18, 19], which has been proven to be a potential biomarker for cancer diagnosis [20]. Therefore, achieving QPI through lensless fiber endomicroscopes could help cancer diagnosis in the early stage with minimum invasiveness. Nevertheless, due to the optical path difference among different fiber cores in the multi-core fiber (MCF), the phase information gets distorted through the lensless endomicroscope [21]. To address this problem, the far-field amplitude-only speckle transfer (FAST) technique [5, 22] is introduced to correct phase distortion via an additional reference measurement, enabling high-fidelity QPI through the lensless fiber endomicroscope. However, the iterative propagation process and the requirement of additional calibration procedure limits FAST’s imaging speed and further applications.
@@
In recent years, deep learning emerges as a powerful tool in QPI [23, 24, 25, 26]. It first verifies its effectiveness in reconstructing phase and amplitude image through a multi-mode fiber [27, 2], successfully reconstructing the handwritten digits in MNIST dataset [28]. Nevertheless, multi-mode fiber is sensitive to fiber conformation and polarization, hindering its ability to achieve high-resolution phase imaging. In contrast, MCFs are relatively less susceptible to external factors. Although deep learning has proved its potential in phase reconstruction through MCFs [29], the capability of the previously reported methods are limited in reconstructing simple structures like handwritten digits and icons [30, 31]. The performance of deep learning models is contingent upon the complexity of the images being reconstructed. Consequently, the relatively simplistic architecture of the conventional U-Net  [32], which is generally utilized in previous works, is inadequate for the phase reconstruction of complex images through scattering medium like MCFs. Diffusion model [33, 34] is the recently prevailing AI model noted for generating high-quality, complex images. Unlike traditional image-to-image reconstruction methods, the generation procedure of diffusion models involves two key processes: the forward process, where data is gradually transformed into Gaussian noise through a sequence of steps, and the reverse process, where the model learns to reconstruct the original data from this noise. The effectiveness of diffusion model origins from its iterative generation process, which alleviates the generation challenge in each step and useful for tasks where fine details are critical. Hence, diffusion models have been widely implemented in sophisticated tasks including drug design [35], protein design [36] and microscopic image data analysis [37, 38]. While typical diffusion models offer powerful capabilities in unsupervised image generation, applying them in complex phase reconstruction tasks remains challenging. Typically, the unsupervised design of diffusion model hinders phase reconstruction from the speckle images captured at the distal far-field of the MCF. The most recent work incorporates physical process into denoising process of the diffusion model, achieving high-quality holographic reconstruction [39]. However, it relies on specific initialization and well-crafted stepsize to avoid undesired artifacts, which undermines the fidelity of the reconstructed images. Therefore, high-fidelity phase reconstruction of complex images using diffusion models remains challenging.
Figure 1: Lensless multi-core fiber phase imaging based on SpecDiffusion. The phase image is displayed on a spatial light modulator and distorted through MCF, forming a speckle pattern on the carema. SpecDiffusion then reconstructs the phase image from a random initialization, guided by the speckle pattern.
@@
In this paper, we introduce a conditional diffusion model, termed SpecDiffusion, to achieve high-fidelity phase reconstruction from the speckle images captured at the distal far-field of the MCF endomicroscope. As illustrated in Fig. 1, SpecDiffusion implements an iterative framework to break the challenging speckle reconstruction task into several steps. Inspired by the phase retrieval algorithms [40], we use the intensity-only speckle images as a supervisor for the diffusion model in each step to ensure high fidelity phase reconstruction. To verify the effectiveness of the proposed diffusion model, we build an optical setup to generate training dataset consisting images from ImageNet [41] and their corresponding speckles through MCF. Given the complexity and diversity of ImageNet images, this dataset can serve as a benchmark for assessing the phase reconstruction capability of the lensless fiber endomicroscope in various scenarios. The pretrained SpecDiffusion model on this dataset demonstrates high robustness in unseen structures. This is validated by a USAF resolution test chart, where the proposed learning-based endomicroscope demonstrates its ability to resolve details down to several micrometers. Additionally, we explore the practical potential of SpecDiffusion in biomedical applications, achieving cancer tissue image reconstruction with high resolution by transfer learning. To further validate the efficacy of SpecDiffusion’s reconstructed tissue images in cell morphological analysis, a zero-shot cell segmentation task is conducted on these reconstructed tissue images. Notably, SpecDiffusion not only achieves improvements in key image quality metrics, but also closely mirrors the segmentation results obtained from the real phase image. Our method offers promising capabilities for phase reconstruction in complex images captured through lensless MCF endomicroscopes, paving the way for its future clinical applications.
2 Results
2.1 Phase reconstruction through conditional diffusion model
@@
Figure 2: Phase reconstruction via SpecDiffusion. a Training process of SpecDiffusion. Phase label is mixed with Gaussian noise, and SpecDiffusion is trained to predict the imposed noise with the guidance of speckle. b Inference process of SpecDiffusion. With the guidance from input speckle, the randomly-generated initial phase is gradually denoised and transformed towards label phase by SpecDiffusion.
@@
We propose a novel conditional diffusion model, termed SpecDiffusion, for phase reconstruction of complex images. It breaks down reconstruction process into iterative denoising steps, gradually transforming the initial random phase into expected phase image. Here, we introduce speckle captured at the distal far-field of MCF as a strong supervisor at each step of the phase reconstruction process. SpecDiffusion is trained from scratch in a supervised way, ensuring the reconstruction fidelity. As illustrated in Fig. 2a, in the training process, the model incrementally adds Gaussian noise to the ground truth phase images over a series of time steps, deteriorating it to a nearly unrecognizable state. Then SpecDiffusion works as a denoising model, predicting the added noise from the noised image. Different from typical diffusion model, each denoising step is supervised by the speckle images captured at the distal side of the MCF, influencing the trajectory of the noise denoising in a physics-informed way.
@@
The inference process employed by the SpecDiffusion is depicted in Fig. 2b. In this process, the model learns to reconstruct the phase information incident the MCF from the noisy random initial phase by performing a series of learned denoising steps. Each denoising step utilizes not only the current state of the reconstructed phase but also the speckle images to reverse the noise addition accurately. Subsequently, this denoised phase is mixed with a smaller amount of Gaussian noise, serving as the initialization for the next iteration. This denoising step is repeated for $T$ iterations, progressively reconstructing the phase image towards the actual phase. The culmination of this process yields the final phase image, which accurately represents the predicted phase corresponding to the input speckle pattern. A detailed explanation of SpecDiffusion is demonstrated in Materials and Methods section.
Figure 3: Scheme of the optical system for generating the far-field speckle dataset. The real phase images are displayed on the SLM and then projected on the MCF facet at the measurement side. The SLM and the detection camera (CAM2222) are synchronously triggered to record the corresponding paired phase and speckle images. LP1111 and LP2222, linear polarizer; L1111-L6666, achromatic lenses; M1 and M2, mirrors; ID, iris diaphragm; BS, beam spliter; CAM1111, alignment camera; MO1111 and MO2222, microscope objectives.
2.2 Experimental setup
To verify the phase reconstruction capacity of SpecDiffusion on complex images, we build an optical system illustrated in Fig. 3. In this system, the phase image is displayed on a Spatial Light Modulator (SLM), then distorted by MCF and finally captured as a speckle pattern on the camera. The detailed description of the optical setup is provided below.
The setup employs a modulated laser beam (Verdi, Coherent Inc.), which is expanded using a telescope system (L1111-L2222) to ensure uniform illumination on the SLM. The SLM dispalys the phase image, thereby modulating the phase of the laser beam. To remove unwanted diffraction orders, a spatial filtering system (L3333-L4444, ID) is utilized. The phase-modulated beam is then projected onto the proximal facet of the MCF (FIGH-$350$S; Fujikura) at the measurement side, using a microscope system (L5555, MO1111; 10×10$\times$10 × plan achromat objective, $0.25$ NA, Olympus). A part of the incident beam is reflected by the MCF facet at the measurement side and directed onto the alignment camera (CAM1111; uEye LE, IDS), facilitating the precise alignment between the holographic display plane and the fiber facet. Far-field speckle images are captured on the detection camera (CAM2222; uEye CP, IDS) through an additional microscope system (MO2222, L6666, LP2222; 10×10$\times$10 × plan fluo objective, $0.3$ NA, Nikon).
In the experiment, the SLM and detection camera are synchronously triggered, ensuring that the phase image projection onto the SLM and the speckle image capture by the camera occur simultaneously. This synchronization guarantees an accurate correspondence between the displayed phase images and the captured speckles. The separation between the far-field image plane and the MCF facet on the detection side measures $0.5$ mm. Compared to traditional endomicroscope system, our system is more simplified and thus better suits the clinical occasion.
2.3 Phase reconstruction on complex images from ImageNet
Utilizing the optical system above, we construct a tailored dataset for QPI in lensless fiber endomicroscopy, based on ImageNet. The capture of images in this dataset begins with converting input grayscale images into phase images. These phase images are resized to a resolution of $980\times 980$ pixels and subsequently zero-padded to $1920\times 1080$ pixels to accommodate the display resolution of the SLM. Computer-generated holograms are then calculated and displayed on the SLM to achieve a precise holographic representation of the ground truth image. Compared to MNIST and Fashion-MNIST datasets in our previous study [29], images from ImageNet are more complex and detailed, thus providing a more robust benchmark for evaluating phase reconstruction capabilities in real-world scenarios. In the experiment, we collect $100,000$ paired phase and speckle images, allocating $90,000$ pairs for training, and the remaining $10,000$ pairs for testing. Details concerning the training methodology and parameters are described in Materials and Methods section.
To visualize the phase reconstruction performance of SpecDiffusion, label phase images are illustrated in Fig. 4a. They are holographically projected on the facet of MCF. When the modulated light field transmitted through the MCF, its phase gets distorted due to intrinsic optical path differences between cores and the honeycomb artifacts [8], thus forming the speckle patterns in Fig. 4b. Based on the speckle patterns, we compare the performance of the U-Net and SpecDiffusion in Fig. 4c and d. Benefiting from the iterative reconstruction process, SpecDiffusion demonstrates the capacity to reconstruct detailed structures within complex images. In the magnified region of the crab image, U-Net manages to reconstruct only the approximate shape of the pincers. In contrast, SpecDiffusion achieves a much clearer and more detailed reconstruction of the pincers. More reconstructed images are illustrated in supplementary Fig. S1111.
@@ 2.3 Phase reconstruction on complex images from ImageNet
Figure 4: Learning-based phase reconstruction of ImageNet images through lensless fiber endomicroscope. a Ground truth phase images. b Speckle patterns captured at the detection side of the lensless fiber endomicroscope. c and d Reconstructed phase images by U-Net and SpecDiffusion. Scale bars $50~{}\mu m$.
To quantitatively characterize the phase reconstruction fidelity, we employ the mean absolute metric (MAE), peak signal-to-noise (PSNR), structural similarity index measure(SSIM) [42] and 2D correlation coefficient as evaluation metrics. The metric distributions of U-Net and SpecDiffusion are illustrated in supplementary Fig. S1111e-h. The averaged evaluation metrics are summarized in Tab. S1111. Quantitatively, SpecDiffusion outperforms conventional U-Net on all metrics, indicating higher fidelity of the reconstructed phase images. The evaluation on ImageNet underlines SpecDiffusion’s ability to enhance diagnostic imaging by providing clearer, more detailed reconstructions that could significantly benefit medical analysis.
2.4 Characterization of endomicroscopic imaging resolution
@@
Figure 5: Learning-based phase reconstruction of test chart through lensless fiber endomicroscope. a Speckle pattern captured at the detection side of lensless fiber endomicroscope . b Ground truth images of the test chart. c U-Net reconstructed phase image. d SpecDiffusion reconstructed phase image. e Phase reconstruction contrast between U-Net and SpecDiffusion. Scale bars $50~{}\mu m$.
SpecDiffusion’s proficiency in reconstructing complex phase images prompted our further investigation into its resolution capabilities. We transform a resolution chart into computer-generated holograms within the phase range of $[0,0.3\pi]$ and then display these holograms on a SLM. The generated speckle patterns are captured by a camera positioned on the detection side of MCF, serving as the input for reconstructing original phase images.
Although the reconstruction of test chart is quite challenging [43], the training process on ImageNet still offers SpecDiffusion resolution capability, indicating its robust phase reconstruction capacity from speckles. As depicted in Fig. 5d, SpecDiffusion achieves a resolution of $4.10~{}\mu m$, surpassing U-Net’s $10.25~{}\mu m$ resolution (see Fig. S2222). Detailedly, in Fig. 5e, the reconstruction results of SpecDiffusion clearly distinguishes the peak phase across three lines and manifest a marked contrast between peaks and valleys within the phase image. This demonstrates that, by reconstructing the speckle pattern, SpecDiffusion can achieve several micrometer level resolution, providing a high-resolution tool for clinical applications.
2.5 Endomicroscopic phase reconstruction of tissue samples
@@
Figure 6: Learning-based cancer tissue reconstruction through the lensless fiber endomicroscope. a Ground truth phase images. b Speckle patterns from MCF captured at the detection side of the lensless fiber endomicroscope. c and d Reconstructed phase image by U-Net and SpecDiffusion. e MAE, f PSNR, g SSIM and h 2D correlation coefficient distribution evaluated on $1,000$ reconstructed tissue images by U-Net and SpecDiffusion. i MAE, j PSNR, k SSIM and l 2D correlation coefficient evaluated on $1,000$ reconstructed tissue images by U-Net and SpecDiffusion with varying tissue dataset size. Scale bars $50~{}\mu m$.
@@
To further evaluate SpecDiffusion’s potential in medical applications, we demonstrate an additional dataset consisting of human cancer tissue. Typically, deep learning models encounter significant challenges with overfitting, resulting in difficulties when generalizing across various application domains, particularly in the medical field. Nevertheless, the diversity inherent in the ImageNet dataset, combined with SpecDiffusion’s robust reconstruction abilities, fortifies our model against overfitting. Notably, with transfer learning on a small set of tissue, which is readily obtainable in clinical scenarios, SpecDiffusion not only sustains its performance but also exhibits improved efficacy. A reconstructed video demonstrating the denoising process of SpecDiffusion is shown in Visualization $1$.

### v4

# Diffusion-driven lensless fiber endomicroscopic quantitative phase imaging towards digital pathology

@@
Abstract
Lensless fiber endomicroscopy is an emerging tool for minimally invasive in vivo imaging, where quantitative phase imaging can be utilized as a label-free modality to enhance image contrast. Nevertheless, current phase reconstruction techniques in lensless multi-core fiber endomicroscopy are effective for simple structures but face significant challenges with complex tissue imaging, thereby restricting their clinical applicability. We present SpecDiffusion, a speckle-conditioned diffusion model tailored to achieve high-resolution and accurate reconstruction of complex phase images in lensless fiber endomicroscopy. Through an iterative refinement of speckle data, SpecDiffusion effectively reconstructs structural details, enabling high-resolution and high-fidelity quantitative phase imaging. This capability is particularly advantageous for digital pathology applications, such as cell segmentation, where precise and reliable imaging is essential for accurate cancer diagnosis and classification. New perspectives are opened for early and accurate cancer detection using minimal invasive endomicroscopy.
Introduction
@@
Lensless fiber endomicroscopy is an emerging tool in medical imaging and diagnostics, noted for its high imaging resolution and minimum invasiveness [1, 2, 3, 4, 5, 6, 7, 8, 9]. Unlike traditional endoscopes that rely on lenses of the application side to capture images, lensless fiber endomicroscope employs computational imaging techniques to visualize microscopic structures with an ultra-thin probe. Therefore, it provides access for previously inaccessible area and reduces neuronal damage during clinical operation, thereby minimizing patient discomfort and contributing to faster post-operation recovery [10, 11, 12]. For typical lensless fiber endomicroscopes employing wavefront shaping technologies, in vivo confocal fluorescent imaging is usually utilized [13, 14, 15, 16, 17, 18]. However, the potential toxin of fluorescent dyes has prompted the exploration of safer alternatives [15], and approaches are being sought to save the effort and time required for staining tissue [19, 20, 21]. Quantitative phase imaging (QPI) serves as a promising label-free technique, which utilizes the phase information to enhance endomicroscopic imaging capabilities [5, 22, 11]. Moreover, it enables the extraction of critical biophysical properties of biological samples, such as refractive index [23, 24], cell volume [25] and dry mass [26, 27], which has been proven to be a potential biomarker for cancer diagnosis [28]. Therefore, achieving QPI through lensless fiber endomicroscopes could help cancer diagnosis in the early stage with minimum invasiveness. Nevertheless, due to the optical path difference among different fiber cores of the multi-core fiber (MCF), the phase information gets distorted through the lensless endomicroscope [29]. To address this problem, the far-field amplitude-only speckle transfer (FAST) technique [5, 30] is introduced to correct phase distortion via an additional reference measurement, enabling high-fidelity QPI through the lensless fiber endomicroscope. However, the iterative propagation process and the requirement of additional calibration procedure limit FAST’s imaging speed. This constraint significantly hamper its application in real-time medical diagnostics during clinical procedures, thereby hindering the precision of a doctor’s clinical operation.
@@
Recently, deep learning emerges as a powerful tool in QPI [31, 32, 33, 34, 35, 36]. It first verifies its effectiveness in reconstructing phase and amplitude image through a multi-mode fiber [37, 2, 38], successfully reconstructing simple structures like the handwritten digits in MNIST dataset [39]. Nevertheless, multi-mode fiber is sensitive to fiber conformation and polarization, hindering its ability to achieve high-resolution phase imaging. In contrast, MCFs are relatively less susceptible to external factors. Although deep learning has proved its potential in phase reconstruction through MCFs [40], the capability of the previously reported methods are limited in reconstructing simple structures like handwritten digits and icons [41, 42]. The performance of deep learning models is contingent upon the complexity of the images being reconstructed. Consequently, the relatively simplistic architecture of the conventional U-Net  [43], which is generally utilized in previous works, is inadequate for the phase reconstruction of complex images through scattering medium like MCFs. Diffusion model [44, 45, 46] is the recently prevailing AI model noted for generating high-quality, complex images. Unlike traditional image-to-image reconstruction methods, the generation procedure of diffusion models involves two key processes: the forward process, where data is gradually transformed into Gaussian noise through a sequence of steps, and the reverse process, where the model learns to reconstruct the original data from this noise. The effectiveness of diffusion model origins from its iterative generation process, which alleviates the generation challenge in each step and useful for tasks where fine details are critical. Hence, diffusion models have been widely implemented in sophisticated tasks including drug design [47], protein design [48] and microscopic image data analysis [49, 50, 51, 48, 52]. While typical diffusion models offer powerful capabilities in unsupervised image generation, applying them in complex phase reconstruction tasks remains challenging. Typically, the unsupervised design of diffusion model hinders phase reconstruction from the speckle images captured at the detection system. The most recent work incorporates physical process into denoising process of the diffusion model, achieving high-quality holographic reconstruction [53]. However, it relies on specific initialization and well-crafted stepsize to avoid undesired artifacts, which undermines the fidelity of the reconstructed images. Therefore, high-fidelity phase reconstruction of complex images using diffusion models remains challenging.
Figure 1: Illustration demonstrates the working principle of the diffusion-driven lensless fiber endomicroscope. The quantitative phase reconstruction process involves using speckle images captured at the detection system to guide the denoising process of the SpecDiffusion model. The reconstructed phase images are applicable to digital pathology tasks, including cell segmentation, enhancing both accuracy and detail in diagnostic imaging.
@@
In this paper, we introduce a conditional diffusion model, termed SpecDiffusion, to achieve high-fidelity quantitative phase reconstruction in the lensless fiber endomicroscope using speckle images captured at the detection system. As illustrated in Fig. 1, SpecDiffusion implements an iterative framework to break the challenging speckle reconstruction task into several steps. Inspired by the phase retrieval algorithms [54], we use the intensity-only speckle images as a supervisor for the diffusion model in each step to ensure high fidelity phase reconstruction. For the phase retrieval process, SpecDiffusion needs only one intensity image as the reference, and the quality of reconstructed phase image is significantly enhanced through the iterative reconstruction process. To verify the effectiveness of the proposed diffusion model, we build an optical setup to generate training dataset consisting images from ImageNet [55] and their corresponding speckles through MCF. Given the complexity and diversity of ImageNet images, this dataset can serve as a benchmark for assessing the phase reconstruction capability of the lensless fiber endomicroscope in various scenarios. The pretrained SpecDiffusion model on this dataset demonstrates high robustness in unseen structures. This is validated by a USAF resolution test chart, where the proposed diffusion-driven endomicroscope demonstrates its ability to resolve details down to several micrometers. Additionally, we explore the practical potential of SpecDiffusion in biomedical applications, achieving cancer tissue image reconstruction with high resolution by transfer learning. To further validate the efficacy of SpecDiffusion’s reconstructed tissue images in digital pathology, a zero-shot cell segmentation task is conducted on these reconstructed tissue images. Notably, SpecDiffusion not only achieves improvements in key image quality metrics, but also closely mirrors the segmentation results obtained from the real phase image. Our method offers promising capabilities for phase reconstruction in complex images captured through lensless MCF endomicroscopes, paving the way for advanced applications in cancer diagnosis and metabolic studies.
Results
Diffusion-driven quantitative phase imaging through a lensless fiber endomicroscope
@@ Results
Figure 2: Architecture of the speckle-conditioned diffusion model (SpecDiffusion). a Training process of SpecDiffusion. Phase label is mixed with Gaussian noise, and SpecDiffusion is trained to predict the imposed noise with the guidance of speckle. b Inference process of SpecDiffusion. With the guidance from input speckle, the randomly-generated initial phase is gradually denoised and transformed towards label phase by SpecDiffusion.
@@
We propose a conditional diffusion model, termed SpecDiffusion, for phase reconstruction of complex images. It breaks down reconstruction process into iterative denoising steps, gradually transforming the initial random phase into desired phase image. The speckle images captured at the far-field of the MCF are served as the conditioning images that guide the iterative denoising process during phase reconstruction. SpecDiffusion is trained from scratch using a supervised approach, ensuring high reconstruction fidelity. As shown in Fig. 2a, during training, Gaussian noise is incrementally added to the ground truth phase images over several time steps, degrading them to an almost unrecognizable state. SpecDiffusion then operates as a denoising model, predicting and progressively removing the noise. Unlike conventional diffusion models, each denoising step is guided by conditioning speckle images, steering the noise reduction process in a physics-informed manner and enhancing the accuracy of phase reconstruction.
@@
The inference process employed by the SpecDiffusion is depicted in Fig. 2b. In this process, the model learns to reconstruct the phase information incident the MCF from the noisy random initial phase by performing a series of learned denoising steps. Each denoising step utilizes not only the current state of the reconstructed phase but also the speckle images to reverse the noise addition accurately. Subsequently, this denoised phase is mixed with a smaller amount of Gaussian noise, serving as the initialization for the next iteration. This denoising step is repeated for $T$ iterations, progressively reconstructing the phase image towards the actual phase. The culmination of this process yields the final phase image, which accurately represents the predicted phase corresponding to the input speckle pattern. Detailed principles of the SpecDiffusion is explained in Methods section.
@@
Figure 3: Diffusion-driven phase reconstruction of ImageNet images through lensless fiber endomicroscope. a Ground truth phase images. b Speckle patterns captured at the detection side of the lensless fiber endomicroscope. c and d Reconstructed phase images by U-Net and SpecDiffusion. The SSIM value for each reconstructed image with respect to its corresponding phase label is shown. Scale bars $50\leavevmode\nobreak\ \mu m$.
To visualize the phase reconstruction performance of SpecDiffusion, ground truth phase images are illustrated in Fig. 3a. They are holographically projected on the facet of MCF. When the modulated light field transmitted through the MCF, its phase gets distorted due to intrinsic optical path differences between cores and the honeycomb artifacts [8], thus forming the speckle patterns in Fig. 3b. Based on the speckle patterns, we compare the performance of the U-Net and SpecDiffusion in Fig. 3c and d. Benefiting from the iterative reconstruction process, SpecDiffusion demonstrates the capacity to reconstruct detailed structures within complex images. In the magnified region of the crab image, U-Net manages to reconstruct only the approximate shape of the pincers. In contrast, SpecDiffusion achieves a much clearer and more detailed reconstruction of the pincers. Further detailed comparison of the phase reconstruction results are demonstrated in supplementary Fig. S2222. To further quantitatively characterize the phase reconstruction fidelity, we employ the mean absolute metric (MAE), peak signal-to-noise (PSNR), structural similarity index measure (SSIM) [56], 2D correlation coefficient and phase residual standard deviation (PRSTD) as evaluation metrics. The metric distributions of U-Net and SpecDiffusion are illustrated in supplementary Fig. S2222e-h. The averaged evaluation metrics are summarized in Table S1111. Quantitatively, SpecDiffusion outperforms conventional U-Net on all metrics, indicating higher fidelity of the reconstructed phase images. The evaluation on ImageNet underlines SpecDiffusion’s ability to enhance diagnostic imaging by providing clearer, more detailed reconstructions. In medical application, clearer reconstructions result in sharper boundaries between different regions, thereby improving discrimination for cancer diagnosis. Moreover, the detailed reconstructions provide more imaging information retrieved from MCF speckles, which is beneficial for further virtual staining.
Characterization of the resolution through the diffusion-driven fiber endomicroscope
@@
Figure 4: Diffusion-driven phase reconstruction of test chart through the lensless fiber endomicroscope. a Speckle pattern captured at the detection side of lensless fiber endomicroscope . b Ground truth images of the test chart. c U-Net reconstructed phase image. d SpecDiffusion reconstructed phase image. e Phase reconstruction contrast of ground truth, U-Net and SpecDiffusion. Scale bars $50\leavevmode\nobreak\ \mu m$.
The evaluation of natural images is crucial for evaluating the general performance of SpecDiffusion. However, to precisely quantify the resolution of the diffusion-driven fiber endomicroscope, it is essential to employ the standardized USAF 1951 test chart as the evaluation target. Reconstructing the test chart using a deep neural network presents a unique challenge because it is an unseen object with distinct high-frequency features that differ significantly from the natural images used to train such models. Deep neural networks, particularly those trained on natural image datasets, are often biased toward capturing low-frequency, global features, which may result in poor reconstruction of fine, high-frequency details like those found in the test chart [57]. This issue is exacerbated in speckle-based imaging systems, where the random and complex nature of speckle patterns further complicates the process of accurately resolving small, detailed structures. As a result, the USAF test chart is a widely recognized benchmark for evaluating the performance of neural networks in resolving fine details, especially when applied to unseen objects with features that differ from the training data.
In the experiment, phase images of the USAF chart were projected onto the spatial light modulator (SLM), with the corresponding speckle patterns captured by the detection system. The reconstructed images were then analyzed to determine the system’s resolution limit by identifying the smallest resolvable elements on the test chart. As depicted in Fig. 4d, the results demonstrate that the diffusion-driven fiber endomicroscope is capable of resolving bars with a width of $4.10\leavevmode\nobreak\ \mu m$, indicating a precision suitable for complex imaging tasks. In comparison, conventional neural networks like U-Net can hardly achieve high resolution reconstruction in this range. As demonstrated in Fig. S3333, the optimum resolution of the system using the U-Net for phase reconstruction is $10.25\leavevmode\nobreak\ \mu m$. In Fig. 4e, the reconstruction results from SpecDiffusion effectively delineate the peak phase across three lines, exhibiting a pronounced contrast between the peaks and valleys within the phase image. This distinct clarity demonstrates that SpecDiffusion, through the accurate reconstruction of speckle patterns, enables micrometer-level imaging resolution for the lensless fiber endomicroscope. This reveals that the diffusion-driven fiber endomicroscope can be a promising high-resolution imaging probe for biomedical applications.
Diffusion-driven fiber-optic phase reconstruction towards human cancer tissue samples
@@
Figure 5: Diffusion-driven cancer tissue reconstruction through the lensless fiber endomicroscope. a Ground truth phase images. b Speckle patterns from MCF captured at the detection side of the lensless fiber endomicroscope. c and d Reconstructed phase image by U-Net and SpecDiffusion. e MAE, f PSNR, g SSIM and h 2D correlation coefficient distribution evaluated on $1,000$ reconstructed tissue images by U-Net and SpecDiffusion. i MAE, j PSNR, k SSIM and l 2D correlation coefficient evaluated on $1,000$ reconstructed tissue images by U-Net and SpecDiffusion with varying tissue dataset size. The SSIM value for each reconstructed image with respect to its corresponding phase label is shown. Scale bars $50\leavevmode\nobreak\ \mu m$.
@@
To further evaluate SpecDiffusion’s potential in biomedical applications, such as cancer diagnosis, we demonstrate an additional dataset consisting of human cancer tissue. Typically, deep learning models encounter significant challenges with overfitting, resulting in difficulties when generalizing across various application domains, particularly in the medical field. Nevertheless, the diversity inherent in the ImageNet dataset, combined with SpecDiffusion’s robust reconstruction abilities, fortifies our model against overfitting. Notably, with transfer learning on a small set of tissue, which is readily obtainable in clinical scenarios, SpecDiffusion not only sustains its performance but also exhibits improved efficacy. A reconstructed video demonstrating the denoising process of SpecDiffusion is shown in Supplementary Video $1$.